OpenStudy (anonymous):

Based on the graph of the following system of equations, determine the number of solutions. 4x - 3y = 8 9y = 12x - 24 Answer Infinitely many solutions Exactly two solutions No solutions Exactly one solution

OpenStudy (eyust707):

did they give you a graph?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

4x-3y=8 12x-9y=-24=>4x-3y=-8 no solution since they are contradicting to each other